Tutor I was wondering if you could help me with these two problems?
1.) Sqrt of 50 (use property 1 to simplify each radical expression. Assume all variables represent positive real numbers.) I got 5 sqrt of 10 is this correct?
2.) sqrt of 72x^3 is the answer 3x sqrt of x?
Thank you for helping me with these problems I appreciate the help.

Answer by stanbon(48568)

1.) Sqrt of 50 (use property 1 to simplify each radical expression. Assume all variables represent positive real numbers.) I got 5 sqrt of 10 is this correct?
sqrt(50) = sqrt(25)*sqrt(2) = 5sqrt(2)
=====================



2.) sqrt of 72x^3 is the answer 3x sqrt of x?
sqrt(72x^3) = sqrt(36x^2)*sqrt(2x)= 6xsqrt(2x)
==================
Cheers,
Stan H.